Supplemental Operations Report 29Apr2023


Supplemental Operations Report – April 29th]

Supplemental Operations Report 29Apr2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation.

Power system:
Solar: nominalSOC Last 24 hours: Max 100%; Min 35 %; Avg 70.3%.
VDC Last 24 hours: Max 58.68V; Min 43.49; Avg 50.90V.Generator run time: about 12am – 6am. Not used on April 28th. Generator hours 3774.7.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 63%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 70%

Water (Static Tank) – 545 gallons
Water in GreenHab – 190 gallons
Water (Outpost tank) – 390 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes.
Hours: 191.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged. All corroded battery wires were replaced. Spare keys were separated from the main ones. All keys are labeled and spare keys placed into the Director’s trailer.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for water runs and supplies pick up.

Crew Car used and why, where? Yes. To Grand Junction by crew. To Hanksville for refill.

General notes and comments: Nominal.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Headset #2 was replaced due to absence of the Ear Pad Cushion.

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Nominal. Hab heater was turned off for tonight.

Summary of GreenHab Operations: Nominal, heater turned off.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation .

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Old wood from the clothesline deck was burned.

Intern’s trailer – heater off.
All keys were labeled.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site